A group of students decided to collect as many rupees for each member of the group as is the number of members. If the total collection amounts to Rs.5929, the number of members in the group is:

Answer by god2012(113)

Solution:
Let 'x' be the total number of members
Total amounted collected from each person = Rs. 'x' .
Therefore , total amount collected from 'x' members = x * x = x^2 = 5929
Taking square root on both sides , we get
x = 77 and x = Rs.77